End of A of a uniform rod AB of weight W and length a is hinged freely to a vertical wall B is connected to C on the wall by a light inextensible string of length a c is at b distant above A show the reaction of hinge at A is w/4b √[a²+8b²]


1
Expert's answer
2021-01-25T13:58:18-0500

Solution:


About the hinge, the torque due to the tension in the string is equal to the torque due to the weight:

T*sin(2Θ)*a = W*("\tfrac{a}{2}")*cosΘ

substitute sin(2Θ) = 2sinΘcosΘ and cancel the a*cosΘ terms:

T*2*sinΘ = "\dfrac{W}{2}"

T = "\dfrac{W}{4sin\theta}"


But sinΘ = "\tfrac{b}{2a}" , so


T = "\dfrac{W}{4(\frac{b}{2a})}=\dfrac{aW}{2b}"

which has vertical component

Ty = T*sinΘ = "(\dfrac{aW}{2b})*(\dfrac{b}{2a})=\dfrac{W}{4}"

and horizontal component

Tx = T*cosΘ = "(\dfrac{aW}{2b})"*"\dfrac{√(a² - b²/4)}{a}" = "(\dfrac{W}{2b})" *"\u221a(a\u00b2 - b\u00b2\/4)"

That makes the reaction components at the hinge

Rx = Tx

and Ry = W - "\dfrac{W}{4}" = "\dfrac{3W}{4}".

Then

R = √["(\dfrac{W}{2b})"²*(a² - "\frac{b^2}{4}") + ("\frac{3W}{4}")²]


R = "(\dfrac{W}{2b})" *√[a² - "\frac{b^2}{4}" + "\frac{9b^2}{4}" ]


R = "(\dfrac{W}{2b})" *√[a² + 2b²] ◄


I have been over this several ways and believe that the "target" you provide is wrong.